OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y = -3x^2 + 14x + 8

myininaya (myininaya):

The vertex of a parabola can be found from this equation y=ax^2+bx+c pretty easily by using the formula for the vertex (-b/(2a), f(-b/(2a)) If a<0, then the parabola is open downwards and has a max. If a>0, then the parabola is open upwards and has a min.

OpenStudy (anonymous):

Okay but I still dont understand how to solve it

myininaya (myininaya):

y=ax^2+bx+c y=-3x^2+14x+8 what is a,b,c?

OpenStudy (anonymous):

I figured out the answer its (14/6,101/3) and 101/3 is the maxium

myininaya (myininaya):

ok the x-coordinate is right... and the y-coordinate is...

myininaya (myininaya):

I got something different for the y-coordinate.

myininaya (myininaya):

-3(14/6)^2 + 14(14/6) + 8

OpenStudy (anonymous):

Oh what was it

myininaya (myininaya):

73/3

myininaya (myininaya):

I gave you a medal for the finding where the max occurred and the actual max is f(-b/2a) which is 73/3 Good job

OpenStudy (anonymous):

Thanks. Goodbye